OpenStudy (anonymous):

Solve for x in the proportion the quantity x plus 1 over 8 equals the quantity 2 times x minus 5 over 2

OpenStudy (haseeb96):

x+1/8 =2(x-5)/2 Now just multiply them and show your answer to me.

OpenStudy (anonymous):

x + 1 ) / 8 = ( 2x - 5 ) / 2 2x + 2 = 16 x - 40 42 = 14 x x = 3